OpenStudy (anonymous):

Please help me!! explain in short sentence 80 kph in 12 minutes, 1 1/2 at 50 kph

OpenStudy (jon123):

Can you expand your question? 1 1/2 what?

OpenStudy (anonymous):

1 and a half at 50 kilometres per hour how long will it take to travel these distances at these speeds

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

please help me anyone!!!!!!!!!!!!!!!!!!

OpenStudy (jon123):

one and a half kilometers at 50 kilometers per hour? If you can go 50 kilometers in one hour, how many hours does it take you to go 100 kilometers? It would take two hours. We find this by dividing the distance by the speed. 100/50 = 2 Likewise, to go 1.5 kilometers, we divide 1.5/50 to get the answer.

OpenStudy (jon123):

For the other part, 80 kilometers in 12 minutes, let's find how many hours 12 minutes is. 60 minutes to an hour, so 12 minutes would be (12/60). Now we want to find the distance? speed*time=distance 80 kph * (12/60) = distance traveled.

OpenStudy (anonymous):

80 multiply by what??

OpenStudy (jon123):

80 kph * time = distance In this case, time is 12 minutes, but since 80 kph uses HOURS, we need to convert 12 minutes to HOURS to be consistent. (12/60) HOURS = 12 minutes

OpenStudy (anonymous):

okay :L

OpenStudy (anonymous):

thanks